# NEAT: Null-Space Constrained Model Editing in A Transfromer

### v2

@@ Abstract
Large language models (LLMs) often exhibit hallucinations due to incorrect or outdated knowledge. Hence, model editing methods have emerged to enable targeted knowledge updates. To achieve this, a prevailing paradigm is the locating-then-editing approach, which first locates influential parameters and then edits them by introducing a perturbation. While effective, current studies have demonstrated that this perturbation inevitably disrupt the originally preserved knowledge within LLMs, especially in sequential editing scenarios. To address this, we introduce AlphaEdit, a novel solution that projects perturbation onto the null space of the preserved knowledge before applying it to the parameters. We theoretically prove that this projection ensures the output of post-edited LLMs remains unchanged when queried about the preserved knowledge, thereby mitigating the issue of disruption. Extensive experiments on various LLMs, including LLaMA3, GPT2-XL, and GPT-J, show that AlphaEdit boosts the performance of most locating-then-editing methods by an average of $36.4\%$ with a single line of additional code for projection solely. Our code is available at: https://github.com/jianghoucheng/AlphaEdit.
@@ 1 Introduction
Large language models (LLMs) have demonstrated the capability to store extensive knowledge during pre-training and recall it during inference (Brown et al., 2020; Petroni et al., 2019; Roberts et al., 2020; Radford et al., 2019). Despite this, they frequently exhibit hallucinations, producing incorrect or outdated information (Cao et al., 2021; Mitchell et al., 2022a). While fine-tuning with updated knowledge offers a straightforward solution, it is often prohibitively time-consuming (Mitchell et al., 2022b). In sight of this, model editing methods have emerged, enabling updating the target knowledge while preserving other knowledge (Meng et al., 2022). Broadly, model editing approaches fall into two categories: (1) parameter-modifying methods, which directly adjust a small subset of parameters (Zhu et al., 2020; Mitchell et al., 2022a; Meng et al., 2022), and (2) parameter-preserving methods that integrate additional modules without altering the original parameters (Mitchell et al., 2022b; Zheng et al., 2023a; Huang et al., 2023; Yu et al., 2024; Hartvigsen et al., 2023).
@@
In this paper, we aim to explore the parameter-modifying methods for model editing. Concretely, current parameter-modifying methods typically follow the locate-then-edit paradigm (Meng et al., 2022; 2023). The basic idea is to first locate influential parameters ${\bm{W}}$ through causal tracing, and then edit them by introducing a perturbation $\bm{\Delta}$ (Meng et al., 2023). The common objective for solving $\bm{\Delta}$ is to minimize the output error on the to-be-updated knowledge, denoted as $e_{1}$. Additionally, the output error on the to-be-preserved knowledge, $e_{0}$, is typically incorporated into the objective function, acting as a constraint to ensure the model’s accuracy on the preserved knowledge.
@@
Figure 1: Comparison between the current methods and AlphaEdit. (a) and (d) exhibit the objectives, where $\lambda$ denotes the coefficient; (b) and (e) show the distributions of hidden representations after dimensionality reduction within the pre-edited and post-edited LLaMA3, respectively; (c) depicts the output of the post-edited LLaMA3. Best viewed in color.
@@
Figure 2: Performance of various model editing methods on LLaMA3 (8B). Results with asterisks in the superscript are from the ZsRE dataset. SST, RTE, and CoLA demonstrate the general capabilities of the post-edited LLMs. The experiments are conducted with 2000 edited samples for sequential editing. Detailed settings are provided in Section 4. Best viewed in color.
@@
Despite their success, the current paradigm faces a critical limitation: it struggles to maintain a balance between knowledge-update error $e_{1}$ and knowledge-preservation error $e_{0}$. Specifically, to prioritize the success of update, prior studies focus more on minimizing $e_{1}$ by assigning a larger weight, while taking insufficient control over $e_{0}$. This could make the LLMs after editing (i.e., the post-edited LLMs) overfit to the updated knowledge. This overfitting would introduce a distribution shift of the hidden representations within LLMs. Figure 2 (b) showcases this shift, where the hidden representations in the post-edited LLaMA-3 (8B) [^1] https://llama.meta.com/lama3/ diverge from their distribution in the original LLM (i.e., the pre-edited LLM). Worse still, in sequential editing scenarios (Gupta & Anumanchipalli, 2024) where the LLM undergoes multiple sequential edits, the accumulation of overfitting gradually erodes the model’s ability to preserve knowledge and generate coherent sentences, eventually leading to model forgetting and model collapse, as depicted in Figure 2 (c).
@@
To validate the effectiveness of our method, we conducted extensive experiments on multiple representative LLMs, such as GPT-2 XL (Radford et al., 2019) and LLaMA-3 (8B), in sequential editing scenarios. The results show that, compared to the best-performing baseline methods, AlphaEdit can achieve over a 36.7% performance improvement on average by adding just one line of code to the conventional model editing method, MEMIT (Meng et al., 2023), as illustrated in Figure 2. Furthermore, we empirically verified that this simple optimization can be easily integrated into most existing model editing methods (Ma et al., 2024; Gu et al., 2024), functioning as a plug-and-play enhancement that significantly boosts their performance.
@@ 2.1 Autoregressive Language Model
This interpretation has inspired most model editing methods to modify the FFN layers for knowledge updates (Meng et al., 2023; Li et al., 2024; Ma et al., 2024). For simplicity, we use ${\bm{W}}$ to refer to ${\bm{W}}_{\text{out}}^{l}$ in the following sections.
@@ 2.2 Model Editing in LLMs
Model editing aims to update knowledge stored in LLMs through a single edit or multiple edits (i.e., sequential editing). Each edit modifies the model parameters ${\bm{W}}$ by adding a perturbation $\bm{\Delta}$ in locate-then-edit paradigm. Specifically, suppose each edit needs to update $u$ new pieces of knowledge in the form of $(s,r,o)$. The perturbed ${\bm{W}}$ is expected to associate $u$ new ${\bm{k}}$-${\bm{v}}$ pairs, where ${\bm{k}}$ and ${\bm{v}}$ encode $(s,r)$ and $(o)$ of the new knowledge, respectively. Let ${\bm{W}}\in\mathbb{R}^{d_{1}\times d_{0}}$, where $d_{0}$ and $d_{1}$ represent the dimensions of the FFN’s intermediate and output layers. Then, we can stack these keys and values into matrices following:
@@ 3.1 Null Space
In the context of model editing, if the perturbation $\bm{\Delta}$ is projected into the null space of ${\bm{K}}_{0}$ (i.e., $\bm{\Delta}{\bm{K}}_{0}=\bm{0}$), adding it to the model parameters ${\bm{W}}$ results in:
$({\bm{W}}+\bm{\Delta}){\bm{K}}_{0}={\bm{W}}{\bm{K}}_{0}={\bm{V}}_{0}.$ (7)
@@ 3.2 Null Space Projecting
As introduced at the end of Section 2.2, the matrix ${\bm{K}}_{0}\in\mathbb{R}^{d_{0}\times 100,000}$ has a high dimensionality with $100,000$ columns. Hence, directly projecting the given perturbation $\bm{\Delta}$ onto the null space of ${\bm{K}}_{0}$ presents significant computational and storage challenges. In sight of this, we adopt the null space of the non-central covariance matrix ${\bm{K}}_{0}{\bm{K}}_{0}^{T}\in\mathbb{R}^{d_{0}\times d_{0}}$ as a substitute to reduce computational complexity, as $d_{0}$ is typically much smaller than $100,000$. This matrix’s null space is consistent with that of ${\bm{K}}_{0}$ (please see Appendix B.2 for detailed proof).
@@ 4.1 Experimental Setup
Base LLMs & Baseline Methods. Our experiments are conducted on three LLMs: GPT2-XL (1.5B), GPT-J (6B), and Llama3 (8B). We compare our method against several model editing baselines, including Fine-Tuning (FT) (Zhu et al., 2020), MEND (Mitchell et al., 2022a), ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), PRUNE (Ma et al., 2024), and RECT (Gu et al., 2024).
Datasets & Evaluation Metrics. We evaluate AlphaEdit using two widely adopted benchmarks: the Counterfact dataset (Meng et al., 2022) and the ZsRE dataset (Levy et al., 2017). In line with prior works (Mitchell et al., 2022a; Meng et al., 2022; 2023), for the Counterfact dataset, we employ Efficacy (efficiency success), Generalization (paraphrase success), Specificity (neighborhood success), Fluency (generation entropy), and Consistency (reference score) as evaluation metrics. For the ZsRE dataset, Efficacy, Generalization, and Specificity are employed.
@@ 4.2 Performance on Knowledge Update and Preservation (RQ1)
  • Obs 2: AlphaEdit enhances text generation fluency and coherence. In addition to editing capabilities, AlphaEdit also exhibits substantial improvements in Fluency and Coherence. For instance, on GPT2-XL, AlphaEdit achieves an $18.33\%$ improvement over the strongest baseline. While other methods often degrade text generation quality to complete the sequential editing task, AlphaEdit preserves both the knowledge and the ability to generate coherent and fluent text.
@@ 4.3 General Capability Tests (RQ2)
    CoLA (Corpus of Linguistic Acceptability) (Warstadt et al., 2019) is a single-sentence classification task, where sentences from linguistic theory books and journals are annotated as either grammatically acceptable or unacceptable.
@@
  • Obs 3: AlphaEdit sustains the general capability of post-edited LLMs even after extensive editing. Specifically, AlphaEdit maintains the original model performance across all metrics, even after editing 3,000 samples. Remarkably, AlphaEdit preserves 98.48% of the model’s general capability on average, demonstrating that the null-space projection not only safeguards the preserved knowledge but also protects the general capability learned from this knowledge’s corpus.
@@ 4.4 Hidden Representations Analysis (RQ3)
Figure 5: The distribution of hidden representations of pre-edited and post-edited LLMs after dimensionality reduction. The top and right curve graphs display the marginal distributions for two reduced dimensions. The dashed lines represent the $0.95$ confidence intervals, where AlphaEdit consistently exhibits minimal shift. Best viewed in color.
@@
  • Obs 5: AlphaEdit maintains consistency in hidden representations after editing. Specifically, the hidden representations within LLMs edited using AlphaEdit remain consistent with the original distribution across all three base models. This stability indicates that AlphaEdit effectively mitigates overfitting, which in turn explains why LLMs edited with AlphaEdit exhibit superior knowledge preservation and generalization capabilities.
@@ 4.5 Performance Improvements of Baseline Methods (RQ4)
We conclude by evaluating whether integrating AlphaEdit’s projection strategy can comprehensively enhance current editing methods. To achieve this, we add one line of code from AlphaEdit (the code for projection) to baselines and measure their performance before and after the addition. Following previous works (Meng et al., 2023; Gu et al., 2024), we analyze (1) the average performance across all editing samples, (2) the performance on editing samples involving specific semantics and (3) the general capability of post-edited LLMs. Results are shown in Figure 6, 7 (a), and 7 (b), respectively. Detailed experimental settings can be found in Appendix C.4. The results show that:
@@
Figure 7: Performance comparison before and after adding the projection code in AlphaEdit to the baselines. (a) Performance of editing methods on samples involving specific semantics, where asterisk denotes the versions added the projection. (b) Comparison of general capabilities for PRUNE and RECT before and after adding the projection, where the underlined method represents the results of RECT. Best viewed in color.
  • Obs 7: AlphaEdit seamlessly integrates with other model editing methods, significantly boosting their overall performance. Specifically, in terms of editing capability, the optimized baseline methods show an average improvement of $28.24\%$, while the general capability improves by $42.65\%$. These results underscore the substantial potential and broad applicability of the null-space projection approach in enhancing model editing approaches.
@@ 5 Related Work
Parameter-modifying Model Editing. This approach typically employs meta-learning or locating-then-editing strategies (Ma et al., 2024; Gu et al., 2024; Zhang et al., 2024b) to conduct editing. Meta-learning, as implemented by KE (Cao et al., 2021) and MEND (Mitchell et al., 2022a), involves adapting model parameters through a hypernetwork. Locate-then-edit strategies, exemplified by ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), prioritize pinpointing the knowledge’s storage location before making targeted edits. Additionally, GLAME Zhang et al. (2024a) enhances current editing methods by leveraging knowledge graphs to facilitate the editing of related knowledge.
Parameter-preserving Model Editing. This line utilizes extra trainable parameters to maintain necessary memories (Wang et al., 2024). SERAC (Mitchell et al., 2022b) constructs an ancillary counterfactual model, utilizing a classifier to determine response applicability. T-Patcher (Huang et al., 2023) integrates additional neurons to rectify each misclassified output. Further innovations include MemPrompt (Madaan et al., 2022) and IKE (Zheng et al., 2023b), which enhance the model’s in-context learning capabilities by introducing modified facts and memory-retrieved demonstrations as prompts. MELO (Yu et al., 2024) dynamically engages specific LoRA blocks, indexed within an internal vector database, to modify model behavior. Conversely, GRACE (Hartvigsen et al., 2023) adopts a unique approach by maintaining a dynamically updated codebook.
Projection-based Continual Learning. It leverages projection to reduce interference between new and old knowledge, which is conceptually similar to model editing. Specifically, OWM (Zeng et al., 2019) directs parameter updates orthogonally to the space of the input data from previous tasks. OGD (Farajtabar et al., 2020) preserves the direction of old gradients and updates along the current gradient direction that is orthogonal to them. Orthog-Subspace (Chaudhry et al., 2020) learns tasks in different (low-rank) vector subspaces to minimize interference. Furthermore, GPM (Saha et al., 2021) identifies the bases of these subspaces by analyzing network activations. TRGP (Lin et al., 2022) introduces “trust regions” to select relevant old tasks for new ones through gradient projections.

### v3

# AlphaEdit: Null-Space Constrained Knowledge Editing for Language Models

@@ Abstract
Large language models (LLMs) often exhibit hallucinations, producing incorrect or outdated knowledge. Hence, model editing methods have emerged to enable targeted knowledge updates. To achieve this, a prevailing paradigm is the locating-then-editing approach, which first locates influential parameters and then edits them by introducing a perturbation. While effective, current studies have demonstrated that this perturbation inevitably disrupt the originally preserved knowledge within LLMs, especially in sequential editing scenarios. To address this, we introduce AlphaEdit, a novel solution that projects perturbation onto the null space of the preserved knowledge before applying it to the parameters. We theoretically prove that this projection ensures the output of post-edited LLMs remains unchanged when queried about the preserved knowledge, thereby mitigating the issue of disruption. Extensive experiments on various LLMs, including LLaMA3, GPT2-XL, and GPT-J, show that AlphaEdit boosts the performance of most locating-then-editing methods by an average of $36.7\%$ with a single line of additional code for projection solely. Our code is available at: https://github.com/jianghoucheng/AlphaEdit.
@@ 1 Introduction
Large language models (LLMs) have demonstrated the capability to store extensive knowledge during pre-training and recall it during inference (Brown et al., 2020; Petroni et al., 2019; Roberts et al., 2020; Liu et al., 2024). Despite this, they frequently exhibit hallucinations, producing incorrect or outdated information (Cao et al., 2021; Mitchell et al., 2022a). While fine-tuning with updated knowledge offers a straightforward solution, it is often prohibitively time-consuming (Mitchell et al., 2022b). In sight of this, model editing methods have emerged, enabling updating the target knowledge while preserving other knowledge (Yao et al., 2023; Gupta et al., 2024). Broadly, model editing approaches fall into two categories: (1) parameter-modifying methods, which directly adjust a small subset of parameters (Meng et al., 2023; Jiang et al., 2025), and (2) parameter-preserving methods that integrate additional modules without altering the original parameters (Huang et al., 2023; Yu et al., 2024; Hartvigsen et al., 2023; Zheng et al., 2023).
@@
In this paper, we aim to explore the parameter-modifying methods for model editing. Concretely, current parameter-modifying methods typically follow the locate-then-edit paradigm (Meng et al., 2022). The basic idea is to first locate influential parameters ${\bm{W}}$ through causal tracing, and then edit them by introducing a perturbation $\bm{\Delta}$ (Li et al., 2024b). The common objective for solving $\bm{\Delta}$ is to minimize the output error on the to-be-updated knowledge, denoted as $e_{1}$. Additionally, the output error on the to-be-preserved knowledge, $e_{0}$, is typically incorporated into the objective function, acting as a constraint to ensure the model’s accuracy on the preserved knowledge.
@@
Figure 1: Comparison between the current methods and AlphaEdit. (a) and (d) exhibit the objectives, where $\lambda$ is the coefficient to keep balance between $e_{0}$ and $e_{1}$ in the objective; (b) and (e) show the distributions of hidden representations after dimensionality reduction within the pre-edited and post-edited LLaMA3, respectively; (c) depicts the output of the post-edited LLaMA3. Best viewed in color.
@@
Figure 2: Performance of various model editing methods on LLaMA3 (8B). Results with asterisks in the superscript are from the ZsRE dataset. SST, RTE, and CoLA demonstrate the general capabilities of the post-edited LLMs. The experiments are conducted with 2000 edited samples for sequential editing. Detailed settings and results are provided in Section 4.2 and Table 1, respectively. Best viewed in color.
@@
Despite their success, the current paradigm faces a critical limitation: it struggles to maintain a balance between knowledge-update error $e_{1}$ and knowledge-preservation error $e_{0}$. Specifically, to prioritize the success of update, prior studies focus more on minimizing $e_{1}$ by assigning a larger weight, while taking insufficient control over $e_{0}$. This could make the LLMs after editing (i.e., the post-edited LLMs) overfit to the updated knowledge. This overfitting would introduce a distribution shift of the hidden representations within LLMs. Figure 2 (b) showcases this shift, where the hidden representations in the post-edited LLaMA-3 (8B) (Meta, 2024) diverge from their distribution in the original LLM (i.e., the pre-edited LLM). Worse still, in sequential editing scenarios (Gupta & Anumanchipalli, 2024) where the LLM undergoes multiple sequential edits, the accumulation of overfitting gradually erodes the model’s ability to preserve knowledge and generate coherent sentences, eventually leading to model forgetting and model collapse, as depicted in Figure 2 (c).
@@
To validate the effectiveness of our method, we conducted extensive experiments on multiple representative LLMs, such as GPT-2 XL (Radford et al., 2019) and LLaMA-3 (8B). The results show that, compared to the best-performing baseline methods, AlphaEdit can achieve over a 36.7% performance improvement on average by adding just one line of code to the conventional model editing method, MEMIT (Meng et al., 2023), as illustrated in Figure 2. Furthermore, we empirically verified that this simple optimization can be easily integrated into most existing model editing methods (Ma et al., 2024; Gu et al., 2024), functioning as a plug-and-play enhancement that significantly boosts their performance. This highlights AlphaEdit’s crucial role in efficient knowledge updates for LLMs, enabling broader applications and future advancements in the field.
@@ 2.1 Autoregressive Language Model
This interpretation has inspired most model editing methods to modify the FFN layers for knowledge updates (Hase et al., 2023; Li et al., 2024a; Hu et al., 2024). For simplicity, we use ${\bm{W}}$ to refer to ${\bm{W}}_{\text{out}}^{l}$ in the following sections.
@@ 2.2 Model Editing in LLMs
Model editing aims to update knowledge stored in LLMs through a single edit or multiple edits (i.e., sequential editing). Each edit modifies the model parameters ${\bm{W}}$ by adding a perturbation $\bm{\Delta}$ in locate-then-edit paradigm. Specifically, suppose each edit needs to update $u$ pieces of knowledge in the form of $(s,r,o)$. The perturbed ${\bm{W}}$ is expected to associate $u$ new ${\bm{k}}$-${\bm{v}}$ pairs, where ${\bm{k}}$ and ${\bm{v}}$ encode $(s,r)$ and $(o)$ of the new knowledge, respectively. Let ${\bm{W}}\in\mathbb{R}^{d_{1}\times d_{0}}$, where $d_{0}$ and $d_{1}$ represent the dimensions of the FFN’s intermediate and output layers. Then, we can stack these keys and values into matrices following:
@@ 3.1 Null Space
In the context of model editing, if the perturbation $\bm{\Delta}$ is projected into the null space of ${\bm{K}}_{0}$ (i.e., ${\bm{\Delta}^{\prime}}{\bm{K}}_{0}=\bm{0}$, where $\bm{\Delta}\textquoteright$ ’ denotes the projected perturbation), adding it to the parameters ${\bm{W}}$ results in:
$({\bm{W}}+{\bm{\Delta}\textquoteright}){\bm{K}}_{0}={\bm{W}}{\bm{K}}_{0}={\bm{V}}_{0}.$ (7)
@@ 3.2 Null Space Projecting
As introduced at the end of Section 2.2, the matrix ${\bm{K}}_{0}\in\mathbb{R}^{d_{0}\times 100,000}$ has a high dimensionality with $100,000$ columns. Hence, directly projecting the given perturbation $\bm{\Delta}$ onto the null space of ${\bm{K}}_{0}$ presents significant computational and storage challenges. In sight of this, we adopt the null space of the non-central covariance matrix ${\bm{K}}_{0}{\bm{K}}_{0}^{T}\in\mathbb{R}^{d_{0}\times d_{0}}$ as a substitute to reduce computational complexity, as $d_{0}$ is typically much smaller than $100,000$. This matrix’s null space is equal to that of ${\bm{K}}_{0}$ (please see Appendix B.2 for detailed proof).
@@ 4.1 Experimental Setup
Base LLMs & Baseline Methods. Our experiments are conducted on three LLMs: GPT2-XL (1.5B), GPT-J (6B), and LLaMA3 (8B). We compare our method against several model editing baselines, including Fine-Tuning (FT) (Zhu et al., 2020), MEND (Mitchell et al., 2022a), InstructEdit (Zhang et al., 2024b), ROME (Meng et al., 2022), MEMIT (Meng et al., 2023), PRUNE (Ma et al., 2024), and RECT (Gu et al., 2024). Furthermore, to further validate the generalizability of AlphaEdit, we include three memory-based editing methods (i.e., SERAC (Mitchell et al., 2022b), GRACE (Hartvigsen et al., 2023) and MELO (Yu et al., 2024)) as baselines in Appendix C.5, along with two additional base LLMs: Gemma (Mesnard et al., 2024) and phi-1.5 (Li et al., 2023) in Appendix C.6.
Datasets & Evaluation Metrics. We evaluate AlphaEdit using two widely adopted benchmarks: the Counterfact dataset (Meng et al., 2022) and the ZsRE dataset (Levy et al., 2017). In line with prior works (Meng et al., 2022), we employ Efficacy (efficiency success), Generalization (paraphrase success), Specificity (neighborhood success), Fluency (generation entropy), and Consistency (reference score) as evaluation metrics. In addition, for comprehensive evaluation, Appendix C.7 presents experiments conducted on three additional datasets: LongformEvaluation (Rosati et al., 2024), MQUAKE (Zhong et al., 2023), and KnowEdit (Zhang et al., 2024d). We encourage interested readers to refer to Appendix C.7 for further details.
@@ 4.2 Performance on Knowledge Update and Preservation (RQ1)
  • Obs 2: AlphaEdit enhances text generation fluency and coherence. In addition to editing capabilities, AlphaEdit also exhibits substantial improvements in Fluency and Coherence. For instance, on GPT2-XL, AlphaEdit achieves an $18.33\%$ improvement over the strongest baseline, demonstrating that it can preserve both the knowledge and the ability to generate fluent text.
@@ 4.3 General Capability Tests (RQ2)
    CoLA (Corpus of Linguistic Acceptability) (Warstadt et al., 2019) is a single-sentence classification task, where sentences are annotated as either grammatically acceptable or unacceptable.
@@
  • Obs 3: AlphaEdit sustains the general capability of post-edited LLMs even after extensive editing. Specifically, AlphaEdit maintains the original model performance across all metrics, even after editing 3,000 samples, demonstrating that the null-space projection not only safeguards the preserved knowledge but also protects the general capability learned from this knowledge’s corpus.
@@ 4.4 Hidden Representations Analysis (RQ3)
Figure 5: The distribution of hidden representations of pre-edited and post-edited LLMs after dimensionality reduction. The top and right curve graphs display the marginal distributions for two reduced dimensions, where AlphaEdit consistently exhibits minimal shift. Best viewed in color.
@@
  • Obs 5: AlphaEdit maintains consistency in hidden representations after editing. Specifically, the hidden representations within LLMs edited using AlphaEdit remain consistent with the original distribution across all three base models. This stability indicates that AlphaEdit effectively mitigates overfitting, in turn explaining its superior knowledge preservation and generalization capabilities.
@@ 4.5 Performance Improvements of Baseline Methods (RQ4)
Figure 7: Performance comparison before and after adding the projection code in AlphaEdit to the baselines. (a) Performance of editing methods on samples involving specific semantics, where asterisk denotes the versions added the projection; the vertical and horizontal axis represent the categories of knowledge and the accuracy of LLM responses involving this knowledge, respectively. (b) Comparison of general capabilities for PRUNE and RECT before and after adding the projection, where the underlined method represents the results of RECT. Best viewed in color.
We conclude by evaluating whether integrating AlphaEdit’s projection strategy can comprehensively enhance current editing methods. To achieve this, we add one line of code from AlphaEdit (the code for projection) to baselines and measure their performance before and after the addition. Following previous works (Meng et al., 2023), we analyze (1) the average performance across all editing samples, (2) the performance on editing samples involving specific semantics and (3) the general capability of post-edited LLMs. Results are shown in Figure 6, 7 (a), and 7 (b), respectively. Note that in Figure 7 (a), the y-axis represents the knowledge belonging to different semantic categories. For instance, the labeled “language” indicates the knowledge instances related to language. Detailed experimental settings can be found in Appendix C.4. The results show that:
  • Obs 7: AlphaEdit seamlessly integrates with other model editing methods, significantly boosting their overall performance. The optimized baselines show an average improvement of $28.24\%$ on editing capability and $42.65\%$ on general capability, underscoring the substantial potential and broad applicability of the null-space projection in enhancing model editing methods.
@@ 5 Related Work
Parameter-modifying Model Editing. This approach typically employs meta-learning or locating-then-editing strategies (Zhang et al., 2024d) to conduct editing. Meta-learning, as implemented by KE (Cao et al., 2021) and MEND (Mitchell et al., 2022a), involves adapting model parameters through a hypernetwork. InstructEdit (Zhang et al., 2024b) extends MEND by designing instructions for training on different tasks. Locate-then-edit strategies, exemplified by ROME (Meng et al., 2022) and MEMIT (Meng et al., 2023), prioritize pinpointing the knowledge’s storage location before making targeted edits. GLAME (Zhang et al., 2024a) enhances ROME by leveraging knowledge graphs to facilitate the editing of related knowledge. Recent work has introduced AnyEdit (Jiang et al., 2025), a recursive approach designed to modify knowledge of arbitrary length and format stored within LLMs.
Parameter-preserving Model Editing. This line utilizes additional modules to store to-be-updated knowledge. These modules may include codebooks, neurons, or auxiliary models, as seen in methods like SERAC (Mitchell et al., 2022b), T-Patcher (Huang et al., 2023), GRACE (Hartvigsen et al., 2023), and MELO (Yu et al., 2024). Additionally, MemPrompt (Madaan et al., 2022) and IKE (Zheng et al., 2023) achieve editing by incorporating to-be-updated knowledge into input prompts. More recently, WISE (Wang et al., 2025) innovates prior module designs by introducing dual memory and conflict-free knowledge sharding, overcoming trade-off between reliability and generalization.
Evaluating Knowledge Editing. Recent work has introduced diverse benchmarks to assess the efficacy of model editing. For instance, KnowEdit (Zhang et al., 2024d) provides a unified datasets by collecting different types of knowledge tailored for knowledge insertion, modification, and erasure tasks; LEME (Rosati et al., 2024), CKnowEdit (Fang et al., 2025), and MQuAKE (Zhong et al., 2023) shift attention to long-form, multi-lingual, and multi-hop knowledge, respectively. These benchmarks collectively push for more comprehensive evaluations.
6 Limitations & Future Discussion
While AlphaEdit demonstrates the capability to edit knowledge with minimal performance degradation, we also recognize its limitations. Concretely, its applicability to multi-modal LLMs (OpenAI et al., 2024) and large reasoning models (Liu et al., 2024) remains unexplored. Hence, future research could focus on extending AlphaEdit to a broader range of base LLMs. Furthermore, the superior performance of null-space projection in balancing knowledge updates and preservation suggests its potential for broader applications. Especially, it could enhance specific LLM capabilities — such as safety, mathematics, or biochemistry — without degrading other abilities. These avenues present exciting opportunities to improve both the applicability and scalability of our method.